# Intrinsic Wrapped Gaussian Process Regression Modeling for Manifold-valued Response Variable

### v1

@@ 4.1 Flight trajectory data
The earth is roughly a sphere and can be modelled as a copy of ${\mathcal{S}^{2}}$, and the flight trajectory can be therefore considered as a manifold-valued random curve. In this article the flight trajectory data from the FlightAware website https://www.flightaware.com/live/flight/CES588/history/20240701/2035Z/KJFK/ZSPD. Specifically, the flight trajectory data used in this study comes from the flight tracking and historical data of Cathay Pacific Airways. These flight trajectories of 20 flights from Hong Kong to London in the period of Jul 30, 2024-Aug 10, 2024. The original data includes time, height, speed, longitude and latitude at discrete time points. To obtain smooth curves from the occasionally noisy data, we pre-smoothed the longitude and latitude data using kernel local linear smoothing and then mapped the longitude and latitude trajectories onto a unit sphere $\mathcal{S}^{2}$, as shown in left panel of Figure 4.
Figure 4: Plots of flight trajectory: (Left) The smooth flight trajectory data from Hong Kong to London at Jul 30, 2024-Aug 10, 2024; (Right) The extensional prediction of flight trajectory based on iWGPR and WGPR. The departure point of the flight is Hong Kong, the destination is London, and the last three points are used as test data.
Table 5: The average and standard errors of the root mean square geodesic errors of WGPR, iWGPR and MGPR for flight trajectory data.
Scheme $N$ iWGPR WGPR MGPR
@@
We select the position of the airplane as the response variable, which can be represented on $S^{2}$ using longitude and latitude, with time considered as covariates. We set the takeoff time of the flight as 0 and the landing time as 1. To account for different levels of sparsity, we consider total data points of $N=30$, $50$, and $100$, chosen at equal intervals from the flight trajectory data. Similar to the previous section, we consider two setups: (1) randomly selecting 20% of the $N$ observation points recorded on August 10, 2024 as test data, and the others as training data (Random); (2) the last five points of the $N$ observation points in each trajectory as test data, and the others as training data (Sort). For the Random setup, the procedure is repeated independently 100 times, and for the Sort setup the 20 flight trajectories are computed separately by using the proposed method. These procedures yield the evaluation metrics: mean and standard errors of the root mean square geodesic errors. Geodesic regression estimation is used as the prior BPF, and a diagonal RBF kernel is taken as the prior TSCF. The prediction accuracies of iWGPR, WGPR and MGPR on the flight trajectory data are presented in Table 5. We see that the iWGPR model has the smallest estimation errors, and outperforms both WGPR and MGPR in terms of interpolation and extrapolation.
In addition, we plot the estimated flight trajectory on August 10, 2024, using two methods, WGPR and iWGPR, as shown in the right panel of Figure 4. Under the Sort sampling scheme, the total number of observation points is $N=40$, with the last three points designated as test data. Compared to WGPR, the estimated trajectory from iWGPR is more close to the true one, indicating that our method, iWGPR, is more effective at modelling manifold data because of its intrinsic property.
(a) True DTI dataset
(b) Samples with 25% of DTI dataset
(c) Samples with 50% of DTI dataset
(d) Reconstruction of DTI dataset
@@ 4.2 Diffusion tensor imaging data
We consider another real data set for a patch of estimated voxel-wise DTI tensors from a coronal slice of an HCP subject [39]. The responses reside on the manifold $\mathcal{S}_{++}^{6}$ and have non-positive curvature. This data structure can be fitted using iWGPR method. The predictors are two-dimensional index variables, which take values in +$\mathbb{Z}^{2}_{+}$. The data used in this study can be downloaded from https://drive.google.com/uc?export=download&id=1psVU_kw86gqYLD9TlyLGvFba61dlWayB. A visualization of the response for DTI data is shown in Figure 5(a). The ellipsoid represents the diffusion characteristics of the DTI tensors, with its shape, size, and orientation determined by the eigenvalues and eigenvectors of the tensors. Additionally, the color represents the relative magnitude of the maximum singular value.
Table 6: The root mean square geodesic errors of WGPR and iWGPR for Diffusion tensor imaging data.
Sampling ratio iWGPR WGPR MGPR
@@
We selected 25%, 50%, and 75% of the total samples, evenly spaced, as train data points, while the remaining samples were used as test data points. The results are reported in Figures 5(b) and 5(c). The primary objective of this study is to reconstruct a patch of estimated voxel-wise Diffusion Tensor Imaging (DTI) tensors from a coronal slice of a subject. To support this reconstruction process, we employed geodesic regression as the prior BPF and used a diagonal Radial Basis Function (RBF) kernel as the TSCF in our iWGPR model. Table 6 presents the root mean square geodesic errors from iWGPR, WGPR and MGPR for the DTI data. We see that iWGPR has considerably smaller estimation errors compared to WGPR and MGPR, highlighting its suerior performance in reconstruction accuracy. Moreover, Figure 5(d) shows the estimated three-dimensional ellipsoid visualization using only 25% of the samples, demonstrating that iWGPR successfully reconstructs the tensor field.
@@ 5 Conclusions
Recognizing that directly defining a covariance function in the ambient space may undermine the inherent geometric properties of a manifold, this paper employs the parallel transport operator to create a new covariance function structure with desirable statistical properties. We then develop an intrinsic wrapped Gaussian process regression model to estimate the regression function and depict the uncertainty. For applications, computation of covariance function of a wrapped Gaussian process usually needs an orthogonal coordinate frame for representation. We demonstrate that the posterior distribution of regression function is independent of the chosen orthogonal coordinate frame. Furthermore, we establish the statistical consistencies of the posterior distribution estimator, including the information consistency and posterior consistency. Numerical studies show that iWGPR is effective in analysing data with Riemannian manifolds, providing meaningful uncertainty estimates.
Our method accommodates data with response in manifold space and predictors in Euclidean space. The exploration of more complex spaces, such as metric spaces, is relatively limited in the existing literature. We anticipate that our proposed method would shed light on analyzing data in complex spaces. In future work, we plan to extend the idea to data where the response is in manifold space and predictors are in non-Euclidean space, developing additive models based on iWGPR.
Although the posterior distribution in this paper has explicit expression, it is no longer a WGP, because the posterior mean ${\boldsymbol{\beta}}_{t^{*}}^{*}$ might not be zero. It raises a question of whether when performing GP regression on manifolds, we should consider other extensions that might be more appropriate than the wrapped extension. Based on non-Euclidean structure of manifolds, constructing a normal distribution directly on a Riemannian manifold without using wrapping function (i.e, the Riemannian exponential map) is significantly challenging. But this area is worth a further study.

### v2

@@ 4.1 Flight trajectory data
The earth is roughly a sphere and can be modelled as a copy of ${\mathcal{S}^{2}}$, and the flight trajectory can be therefore considered as a manifold-valued random curve. In this article, the flight trajectory data is from the FlightAware website (https://www.flightaware.com/). Specifically, the flight trajectory data used in this study come from the flight tracking and historical data of Cathay Pacific Airways. These flight trajectories of 20 flights from Hong Kong to London in the period of Jul 30, 2024-Aug 10, 2024. To obtain smooth curves from the occasionally noisy data, we pre-smoothed the longitude and latitude data using kernel local linear smoothing and then mapped the longitude and latitude trajectories onto a unit sphere $\mathcal{S}^{2}$.
We select the position of the airplane as the response variable, which can be represented on $S^{2}$ using longitude and latitude, with time considered as covariates. We set the takeoff time of the flight as 0 and the landing time as 1. To account for different levels of sparsity, we consider total data points of $N=30$, $50$, and $100$, chosen at equal intervals from the flight trajectory data. Similar to the previous section, we consider two setups: (1) randomly selecting 20% of the $N$ observation points recorded on August 10, 2024 as test data, and the others as training data (Random); (2) the last five points of the $N$ observation points in each trajectory as test data, and the others as training data (Sort). For the Random setup, the procedure is repeated independently 100 times, and for the Sort setup the 20 flight trajectories are computed separately by using the proposed method. These procedures yield the evaluation metrics: mean and standard errors of the root mean square geodesic errors. Geodesic regression estimation is used as the prior BPF, and a diagonal RBF kernel is taken as the prior TSCF. The prediction accuracies of iGPR, WGPR and MGPR on the flight trajectory data are presented in Table 3. We see that the iGPR model has the smallest estimation errors, and outperforms both WGPR and MGPR in terms of interpolation and extrapolation.
Table 3: The average and standard errors of the root mean square geodesic errors of WGPR, iGPR and MGPR for flight trajectory data.
Scheme $N$ iGPR WGPR MGPR
@@
In addition, we plot the estimated flight trajectory on 10 August 2024, using two methods, WGPR and iGPR, as shown in Figure of Section 10 of the Supplementary Materials. Under the Sort sampling scheme, the total number of observation points is $N=40$, with the last three points designated as test data. Compared to WGPR, the estimated trajectory from iGPR is closer to the true one, indicating that our method, iGPR, is more effective at modeling manifold data because of its intrinsic property.
@@ 4.2 Diffusion tensor imaging data
We consider another real data set for a patch of estimated voxel-wise DTI tensors from a coronal slice of an HCP subject [17]. The responses reside on the manifold $\mathcal{S}_{++}^{6}$ and have a non-positive curvature. This data structure can be fitted using the iGPR method. The predictors are two-dimensional index variables, which take values in +$\mathbb{Z}^{2}_{+}$. The data used in this study can be downloaded from (https://drive.google.com/uc?export=download&id=1psVU_kw86gqYLD9TlyLGvFba61dlWayB). A visualization of the response for DTI data is shown in Figure 3(a). The ellipsoid represents the diffusion characteristics of the DTI tensors, with its shape, size, and orientation determined by the eigenvalues and eigenvectors of the tensors. Additionally, the color represents the relative magnitude of the maximum singular value.
Table 4: The root mean square geodesic errors of WGPR and iGPR for Diffusion tensor imaging data.
Sampling ratio iGPR WGPR MGPR
@@
(a) True DTI dataset
(b) Reconstruction of DTI dataset
We selected 25%, 50%, and 75% of the total samples, evenly spaced, as train data, while the remaining samples were used as test data. The primary objective of this study is to reconstruct a patch of estimated voxel-wise DTI tensors from a coronal slice of a subject. To support this reconstruction process, we employed geodesic regression as the prior BPF and used a diagonal Radial Basis Function (RBF) kernel as the TSCF in iGPR. Table 4 presents the root mean square geodesic errors from iGPR, WGPR and MGPR for DTI data. We see that iGPR has considerably smaller estimation errors compared to WGPR and MGPR, highlighting its superior performance in reconstruction accuracy. Moreover, Figure 3(b) shows the estimated three-dimensional visualization of the ellipsoid using only 25% of the samples, demonstrating that iGPR successfully reconstructs the tensor field.
@@ 5 Conclusions
Recognizing that directly defining a covariance function in the ambient space may undermine the inherent geometric properties of a manifold, this paper employs the parallel transport operator to create a new structure of the covariance function with desirable statistical properties. We then develop an intrinsic Gaussian process regression model to estimate the regression function and depict the uncertainty. For applications, computation of covariance function of a intrinsic Gaussian process usually needs an orthogonal coordinate frame for representation. We show that the posterior distribution of the regression function is independent of the chosen orthogonal coordinate frame. Furthermore, we establish the statistical consistencies of the posterior distribution estimator, including the information consistency and posterior consistency. Numerical studies show that iGPR is effective in analyzing data with Riemannian manifolds, providing meaningful uncertainty estimates.
Our method accommodates data with response in the manifold space and predictors in Euclidean space. The exploration of more complex spaces, such as metric spaces, is relatively limited in the existing literature. We anticipate that our proposed method would shed light on the analysis of data in complex spaces. In future work, we plan to extend the idea to data where the response is in manifold space and predictors are in non-Euclidean space, developing additive models based on iGPR. Although the posterior distribution in this paper has explicit expression, it is no longer a iGD, because the posterior mean ${\boldsymbol{\beta}}_{t^{*}}^{*}$ might not be zero. It raises a question of whether when performing GP regression on manifolds, we should consider other extensions that might be more appropriate than the extension. Based on non-Euclidean structure of manifolds, constructing a normal distribution directly on a Riemannian manifold without using the Riemannian exponential map is significantly challenging. However, this area is worth further study.